# Learning Confident Classifiers in the Presence of Label Noise

### v2

@@ B.1 Segmentation Datasets
It is the database of handwritten digits that is used as an example of a dataset with artificial noises. This dataset is used to show the performance of the model in cases where ground truth is known. Synthetic annotations are created by applying morphological transformations simulations on assumed ground truth [8]. The types of annotators from the simulation are thin (under-segmentation), thick (over-segmentation), fracture (fracture-segmentation), and good (ground truth). For our purpose, we only use the first three annotations. In addition, Gaussian noise was added to the original images to use them as inputs for the model. There are 60,000 samples for training and 10,000 data samples reserved for testing. In Figure 6, we show different augmentations for the MNIST dataset, artificially creating several annotators.
@@
B.2 Classification Datasets
MNIST.
As described in Appendix B.1, dataset comprise of 60,000 samples for training, and 10,000 data samples reserved for testing. The number of classes in the dataset is 10.
CIFAR-10.
The CIFAR-10 dataset contains 60,000 color images in 10 classifications, with 6000 images in each class. There are 50,000 training and 10,000 test images. The dataset is divided into five training batches and one test batch, each contains 10,000 images. The test batch is a collection of exactly 1,000 data samples randomly selected from each class. The training batches comprise of the remaining images in a random order, however it’s likely that certain training batches may have more images from one class than another.
CIFAR-10N.
CIFAR-10N is prepared by augmenting the training datasets of CIFAR-10 using Amazon Mechanical Turk-collected human-annotated real-world noisy labels. It has the same number of classes and training and test data points as CIFAR-10 dataset.
FMNIST.
B.2.1 Types of Noise
@@ C.1 Segmentation
We tuned relative weight hyperparameters such as learning rate, $\lambda,\gamma$ for segmentation model using automatic tool (wandb.ai) for the results in the main text and ablation study in Section C.2. Then, we picked the best values in terms of the Dice score between predicted and true labels.
@@
Ours (No Confidence) class 0 91.01 90.90 98.47 95.05 95.34 97.39
@@
Ours (No Confidence) class 0 100.00 99.83 99.57 99.65 99.86 99.65
@@
Ours (No Confidence) class 0 99.62 98.96 99.05 99.24 99.35 99.26
@@
C.2 Comparative Analysis of Various Regularizers for Segmentation.
@@
The results of our analysis, which are summarized in Table 9, consider several regularizers. Among these are the CM-Net [56] and the Logdet [29], in addition to other combinations. Our intention was to compare these well-established regularizers against the one we propose in this study.
Notably, the table reveals that our proposed regularizer, which includes an additional term for confident regions, exceeds the performance of the other options by a considerable margin. This superior performance may be attributed to the fact that our proposed regularizer does not impose any additional constraints on the parameters. For instance, it does not enforce diagonal dominance as in the Logdet [29], a requirement that is explicitly stated. Neither does it employ the same constraints as those theoretically proven in CM-Net [56], whose practical implementation in the algorithm and code remains somewhat ambiguous.
@@
Table 9: In this table we present additional ablation results. Average Dice score with standard deviation over 10 last epochs, computed for different combinations of regularizers. Reg-(Y/N) means whether we apply (Y) our new regularizer or not (N). CR-(Y/N) means the same for the confidence region term on the loss function. The symbol “R” represents the count of dilation iterations. In the last two columns, we keep Reg-Y and CR-Y, while varying the number of dilation iterations. The bold results highlight the top performances, while the underlined ones signify the next best outcomes.
@@ C.3 Classiffication
We used a LeNet model as a classifier for our backbone network. For the annotator network, we have a linear layer of size $C\times C$, $C$ denotes the number of classes. This linear layer represents our annotator confusion matrices, and we apply a softmax layer to it to make it a stochastic matrix along a certain dimension. We fine-tuned our model for a combination of learning rates, $\alpha$ = [0.01, 0.001, 0.0001, 0.000001, 0.0016, 0.008, 0.0064, 0.005], and about 50 different lambda values, $\lambda$, for our regularizer hyper-parameter. We started with a very small value of $\lambda$ = 0.001506746, and slowly increased it exponentially (geometric progression) per epoch with a rate, r = 1.18; we trained the model for 70 epochs and used Adam as an optimizer. In addition, the experiments were regulated to assess the performance of the model when the confusion matrix is initialized as an identity. These fine-tunings are done across all 2 different types of noise described in Section 5.2.3 with the respective noise rate that is associated with each of the noise types.
Table 10 shows the performance comparison of our algorithm with the other methods. The results of the our methodology with entropy and information-based regularizers are superior compared to other method, specially for more challenging cases of higher noise rates.
@@
[…]
Figure 15: Test accuracy (%) vs. number of epochs on MNIST.
@@
Figure 15: Test accuracy (%) vs. number of epochs on MNIST.
[…]
@@
For CIFAR-10, we used ResNet-18 as our backbone network for the classifier. The annotator network remains unchanged (still has one linear layer that represents the confusion matrices of class $C\times C$ that are stochastic). For this dataset, we fine-tuned the model for an assortment of learning rates, such as $\alpha$ = [0.001, 0.00064, 0.0016, 0.000001, 0.005, 0.008, 0.0016, 0.00064]. We ran the model for 150 epochs; the hyperparameter $\lambda$ for our regularizer was slowly increased exponentially again with a rate, r= 1.11. However, the starting value this time is $\lambda$= 3.0517578125e-05. We used a standard batch-size, BS=128. We used the standard augmentations of random crop of size 32 $\times$ 32 and horizontal random flipping. These are the standard augmentations that have been used across all the baselines that we have evaluated. The remaining settings remain the same as described in MNIST above.
@@
We kept the same settings of CIFAR-10, such as ResNet-18 model and batch-size of 128 for Fashion-MNIST dataset. The model was again fine-tuned for the same set of hyperparameters. However, the starting value of $\lambda$= 6.103515625e-05, and it was increased exponentially with a rate of r=1.12. We also retained the same set of augmentations that we used in CIFAR-10 dataset.
@@
We kept the same backbone architecture and fine-tune settings as CIFAR-10 in Section C.3.
Refer to caption (a) Refer to caption (b) Refer to caption (c)
@@
C.3.1 Additional experimental results
@@
C.3.2 MNIST Curated Dataset
In addition to symmetric and pairflip noises described in Section 5.2.3, we also consider asymmetric and pairflip with permutation noises. In the latter, the ordered label categories were first permuted randomly and labels of two adjacent categories after permutation were swapped based on a preset ratio. Asymmetric noise is a block matrix transformation, where a portion of original labels are retained and the remainder is uniformly reassigned to closest four categories.
In Figure 19 we demonstrate annotators’ confusion using our algorithm on the curated MNIST dataset that showcases different image styles of Original, Thin and Thick. The strength of the regularizer, $\lambda$=0.01, is increased by the multiplicative scalar m=2 every epoch. Figures 20, 21 and 22 highlights the original and predicted confusion of annotator 1, annotator 2 and annotator 3 using our approach with the regularizer and the non-regularized approach (that is, when $\lambda$=0).
@@ C.3.2 MNIST Curated Dataset
Figure 19: Learned Annotators’ confusion for different image styles using our approach with the regularizer ($\lambda$=0.01, m=2) on MNIST dataset.
Image Original Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 20: Original and Predicted confusion for Annotator 1 using different models: our approach with regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on MNIST dataset.
Image Original Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 21: Original and Predicted confusion for Annotator 2 using different models: our approach with regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on MNIST dataset.
Image Original Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 22: Original and Predicted confusion for Annotator 3 using different models: our approach with regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on MNIST dataset.

### v3

@@ C.2 Segmentation Datasets
It is the database of handwritten digits that is used as an example of a dataset with artificial noises. This dataset is used to show the performance of the model in cases where ground truth is known. Synthetic annotations are created by applying morphological transformation simulations on assumed ground truth [8]. The types of annotators from the simulation are thin (under-segmentation), thick (over-segmentation), fracture (fracture-segmentation), and good (ground truth). For our purpose, we only use the first three annotations. In addition, Gaussian noise was added to the original images to use them as inputs for the model. There are 60,000 samples for training and 10,000 data samples reserved for testing. In Figure 6, we show different augmentations for the MNIST dataset, artificially creating several annotators.
@@
C.2.1 Types of Noise
@@ D.1 Segmentation
We tuned relative weight hyperparameters such as learning rate, $\lambda,\gamma$ for the segmentation model using an automatic tool (wandb.ai) for the results in the main text and ablation study in Section D.2. Then, we picked the best values in terms of the Dice score between predicted and true labels.
@@
Ours (No Conf.) class 0 91.01 90.90 98.47 95.05 95.34 97.39
@@
Ours (No Conf.) class 0 100.00 99.83 99.57 99.65 99.86 99.65
@@
Ours (No Conf.) class 0 99.62 98.96 99.05 99.24 99.35 99.26
@@
D.2 Comparative Analysis of Various Regularizers for Segmentation.
@@
The results of our analysis, which are summarized in Table 8, consider several regularizers. Among these are the CM-Net [54] and the Logdet [28], in addition to other combinations. We intended to compare these well-established regularizers against the one we propose in this study.
Notably, the table reveals that our proposed confidence regularizer, which includes an additional term for confident regions (CR), exceeds the performance of the other options by a considerable margin. This superior performance may be attributed to the fact that our proposed regularizer does not impose any additional constraints on the parameters. For instance, it does not enforce diagonal dominance as in the Logdet [28], a requirement that is explicitly stated. Neither does it employ the same constraints as those theoretically proven in CM-Net [54], whose practical implementation in the algorithm and code remains somewhat ambiguous.
Table 8: In this table we present additional ablation results. Average Dice score with standard deviation over 10 last epochs, computed for different combinations of regularizers. Reg-(Y/N) means whether we apply (Y) our confidence regularizer or not (N). CR-(Y/N) means the same for the confidence region (CR) term on the loss function. The symbol “R” represents the count of dilation iterations. In the last two columns, we keep Reg-Y and CR-Y, while varying the number of dilation iterations. The bold results highlight the top performances, while the underlined ones signify the next best outcomes.
@@ D.3 Classiffication
We used a LeNet model as a classifier for our backbone network. The LeNet5 architecture is a convolutional neural network for image classification, featuring a feature extractor and a classifier. The feature extractor consists of three convolutional layers with 6, 16, and 120 filters, each followed by ReLU activation and average pooling. The output is flattened and passed to the classifier, which has two fully connected layers: the first with 84 nodes and the second with number of classes nodes, ending with a softmax for class probabilities. For the annotator network, we have a linear layer of size $C\times C$, $C$ denotes the number of classes. This linear layer represents our annotator confusion matrices, and we apply a softmax layer to it to make it a stochastic matrix along a certain dimension. We fine-tuned our model for a combination of learning rates, $\alpha$ = [0.01, 0.001, 0.0001, 0.000001, 0.0016, 0.008, 0.0064, 0.005], and about 50 different lambda values, $\lambda$, for our regularizer hyper-parameter. We started with a very small value of $\lambda$ = 0.001506746, and slowly increased it exponentially (geometric progression) per epoch with a rate, r = 1.18; we trained the model for 70 epochs and used Adam as an optimizer. In addition, the experiments were regulated to assess the performance of the model when the confusion matrix was initialized as an identity. These fine-tunings are done across all 2 different types of noise described in Section 2 with the respective noise rate that is associated with each of the noise types.
Tables 10 and 10 show the performance comparison of our algorithm with the other methods. The results of our methodology with entropy and confidence regularizers are superior compared to other methods, especially for more challenging cases of higher noise rates.
@@
Noise rate Ours-Inf Ours-Ent Co-tea Co-tea+ JoCoR Trace CDR
symmetric 20% 90.67 90.79 90.48 88.69 91.88 90.61 88.69
$\pm$ 0.21 $\pm$ 0.19 $\pm$ 0.06 $\pm$ 0.09 $\pm$ 0.06 $\pm$ 0.24 $\pm$ 0.09
symmetric 30% 91.35 90.34 90.36 88.50 91.33 89.64 87.38
$\pm$ 0.25 $\pm$ 0.26 $\pm$ 0.13 $\pm$ 0.09 $\pm$ 0.08 $\pm$ 0.30 $\pm$ 0.09
symmetric 50% 89.51 89.49 89.37 77.96 89.21 88.94 85.36
$\pm$ 0.55 $\pm$ 0.58 $\pm$ 0.08 $\pm$ 0.33 $\pm$ 0.06 $\pm$ 0.53 $\pm$ 0.09
pairflip 20% 90.90 90.77 90.68 89.12 91.37 90.40 90.01
$\pm$ 0.33 $\pm$ 0.29 $\pm$ 0.05 $\pm$ 0.08 $\pm$ 0.09 $\pm$ 0.35 $\pm$ 0.11
pairflip 30% 90.38 90.65 90.11 89.06 89.67 90.33 88.78
$\pm$ 0.41 $\pm$ 0.45 $\pm$ 0.08 $\pm$ 0.10 $\pm$ 0.07 $\pm$ 0.46 $\pm$ 0.14
pairflip 45% 89.37 89.02 78.86 52.61 88.10 89.08 64.63
$\pm$ 0.68 $\pm$ 0.71 $\pm$ 0.11 $\pm$ 0.43 $\pm$ 0.39 $\pm$ 0.72 $\pm$ 0.17
@@
Figure 15: Test accuracy (%) vs. number of epochs on MNIST.
@@
For CIFAR-10, we used ResNet-18 as our backbone network for the classifier. The ResNet18 architecture has 18 learnable layers. It starts with an initial convolutional layer that outputs 64 filters, followed by 4 residual blocks where the number of filters increases from 64 to 512. After global average pooling reduces the feature map to 512 dimensions, a custom fully connected layer maps this to the target classes for classification. The annotator network remains unchanged (still has one linear layer that represents the confusion matrices of class $C\times C$ that are stochastic). For this dataset, we fine-tuned the model for an assortment of learning rates, such as $\alpha$ = [0.001, 0.00064, 0.0016, 0.000001, 0.005, 0.008, 0.0016, 0.00064]. We ran the model for 150 epochs; the hyperparameter $\lambda$ for confidence regularizer was slowly increased exponentially again with a rate, r= 1.11. However, the starting value this time is $\lambda$= 3.0517578125e-05. We used a standard batch size, BS=128. We used the standard augmentations of a random crop of size 32 $\times$ 32 and horizontal random flipping. These are the standard augmentations that have been used across all the baselines that we have evaluated. The remaining settings remain the same as described in MNIST above.
@@
We kept the same settings of CIFAR-10, such as the ResNet-18 model and batch size of 128 for the Fashion-MNIST dataset. The model was again fine-tuned for the same set of hyperparameters. However, the starting value of $\lambda$= 6.103515625e-05, and it was increased exponentially with a rate of r=1.12. We also retained the same set of augmentations that we used in the CIFAR-10 dataset.
@@
We kept the same backbone architecture and fine-tune settings as CIFAR-10 in Section D.3.
@@
D.3.1 Additional experimental results
@@
D.3.2 Curated MNIST Dataset
In addition to symmetric and pairflip noises described in Section 2, we also consider asymmetric and pairflip with permutation noises. In the latter, the ordered label categories were first permuted randomly and labels of two adjacent categories after permutation were swapped based on a preset ratio. Asymmetric noise is a block matrix transformation, where a portion of original labels are retained and the remainder is uniformly reassigned to the closest four categories.
In Figure 19 we demonstrate annotators’ confusion using our algorithm on the Curated MNIST dataset that showcases different image styles of Original, Thin, and Thick. The strength of the confidence regularizer, $\lambda$=0.01, is increased by the multiplicative scalar m=2 every epoch. Figures 20, 21 and 22 highlight the original and predicted confusion of annotator 1, annotator 2, and annotator 3 using our approach with the regularizer and the non-regularized approach (that is, when $\lambda$=0).
@@ D.3.2 Curated MNIST Dataset
Figure 19: Learned Annotators’ confusion for different image styles using our approach with the confidence regularizer ($\lambda$=0.01, m=2) on MNIST dataset.
Image Ground Truth Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 20: Ground truth and predicted confusion for Annotator 1 using different models: our approach with confidence regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on Curated MNIST dataset.
Image Ground Truth Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 21: Grount truth and predicted confusion for Annotator 2 using different models: our approach with confidence regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on Curated MNIST dataset.
Image Ground Truth Our ($\lambda$=0.01, m=2) Our ($\lambda$=0)
@@
Figure 22: Ground truth and predicted confusion for Annotator 3 using different models: our approach with confidence regularizer ($\lambda=0.01$, m=2) and without it ($\lambda=0$) on Curated MNIST dataset.